### v1

# DECIPHERING THE DEFINITION OF ADVERSARIAL ROBUSTNESS FOR POST-HOC OOD DETECTORS

Abstract
Detecting out-of-distribution (OOD) inputs is critical for safely deploying deep learning models in real-world scenarios. In recent years, many OOD detectors have been developed, and even the benchmarking has been standardized, i.e. OpenOOD. The number of post-hoc detectors is growing fast and showing an option to protect a pre-trained classifier against natural distribution shifts, claiming to be ready for real-world scenarios. However, its efficacy in handling adversarial examples has been neglected in the majority of studies. This paper investigates the adversarial robustness of the 16 post-hoc detectors on several evasion attacks and discuss a roadmap towards adversarial defense in OOD detectors.
Machine Learning, ICML, OOD, adversarial examples
1 INTRODUCTION
Adversarial robustness in the context of out-of-distribution (OOD) detection refers to the ability of a detector to correctly identify OOD samples even when they have been adversarial perturbed to evade a deep neural network (DNN). Evasion attacks, e.g. PGD (Madry et al., 2017), which are designed to fool deep learning classifiers, are difficult to spot as an outlier for OOD detectors. To prevent errors in real-world applications, it is crucial to detect OOD cases, not only for natural distribution shifts (Taori et al., 2020) but also for adversarial examples (Sehwag et al., 2019a) without degrading the generalizability of the underlying pre-trained classifier (Yang et al., 2021).
Current standardized benchmarks such as OpenOOD (Zhang et al., 2023) and RoboDepth (Kong et al., 2024) merely focus on natural distribution shifts and corruptions (Hendrycks & Dietterich, 2019). Especially OpenOOD aims to make a fair comparison across methods initially developed for anomaly detection, model uncertainty, open set recognition, and OOD detection methods.
OpenOOD benchmark suite (Yang et al., 2021) evaluates methods on semantic shift (e.g., samples that are semantically different from the training data, representing truly novel or unseen concepts.) (Hendrycks & Gimpel, 2016) and covariate shift (e.g. samples that come from a different distribution than the training data, but still belong to the same semantic categories).
[…]
Detector Venue Detector Architecture Adversarial
Features Logits Probs Robust
SCALE ICLR’24 $\checkmark$ $\checkmark$
NNGUIDE NeurIPS’23 $\checkmark$ $\checkmark$
GEN CVPR’23 $\checkmark$
ASH ICLR’23 $\checkmark$
DICE ECCV’22 $\checkmark$
KNN ICML’22 $\checkmark$
VIM CVPR’22 $\checkmark$ $\checkmark$
KLM ICML22 $\checkmark$
MLS ICML22 $\checkmark$
REACT NeurIPS’21 $\checkmark$
[…]
GRAM ICML’20 $\checkmark$
EBO NeurIPS’20 $\checkmark$
ODIN ICLR’18 $\checkmark$
[…]
MSP ICLR’17 $\checkmark$
Current OOD detection methods, as some listed in Table 1, achieve outstanding results on prominent OOD benchmarks, such as the OpenImage-O (Wang et al., 2022), ImageNet-O (Hendrycks et al., 2021), Texture (Cimpoi et al., 2014), and iNaturalist (Huang & Li, 2021; Van Horn et al., 2018). OOD detection is a very quickly growing field due to the number of methods added to OpenOOD. More specific post-hoc methods with their plug-and-play capabilities on pre-trained classifiers are more flexible and scalable compared to methods that require full retraining on new OOD data (Yang et al., 2022a; Cong & Prakash, 2022). Simple post-hoc methods like KNN (Sun et al., 2022) are highlighted maintaining good performance on toy-datasets (e.g. MNIST (Deng, 2012), CIFAR-10 or CIAR-100 (Krizhevsky et al., 2009), and also show outstanding performance on the more realistic dataset like ImageNet (Deng et al., 2009) according to (Yang et al., 2022a). These experiments do neglect the adversarial robustness of “state-of-the-art” detectors and the real-world capabilities are questionable as past studies had shown (Sehwag et al., 2019a; Song et al., 2020; Chen et al., 2020; Salehi et al., 2021). Adversarial examples remain challenging because they share the same semantics as the training data, but aim to modify the classifier’s output.
In this study, we investigate the adversarial robustness of post-hoc ODD detectors. Our contributions can be summarized as follows:
  • We revise the definition of adversarial OOD regarding post-hoc OOD methods to finally have a common understanding of robust adversarial OOD detection.
  • We examine 16 post-hoc OOD detectors by delving into their current ability to detect adversarial examples — an aspect that has been disregarded.
  • We expand the OpenOOD framework with evasive attacks and provide adversarial OOD datasets: github.com/adverML/AdvOpenOOD.
2 RELATED WORK
2.1 Evasion Attacks Crafting Inliers
The objective of evasion attacks is to generate adversarial examples that will result in the misclassification of inputs at deep learning models (Biggio et al., 2013). It is possible to distinguish between two types of attack: black-box attacks (Zheng et al., 2023), where the classifier is queried, and white-box attacks (Carlini & Wagner, 2017b), where the network is under the attacker’s complete control. A white-box threat model is strictly stronger. They try to find the smallest possible perturbation, often imperceptible to humans, to manipulate the model’s decision boundaries.
More formally, for an input ${\bm{x}}$ with the ground-truth label $y$, an adversarial example ${\bm{x}}^{\prime}$ is crafted by adding small noise $\bm{\delta}$ to ${\bm{x}}$ such that the predictor model loss ${\bm{J}}({\bm{x}}^{\prime},y)$ is maximized. The $L^{p}$ norm of the adversarial noise should be less than a specified value $\epsilon$, i.e., $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, e.g., $\epsilon=8/255$ (Croce et al., 2020), to ensure that the image does not change semantically. The attack method Fast Gradient Sign Method (FGSM) by (Goodfellow et al., 2014) maximizes the loss function in a single step by taking a step towards the sign of the gradient of ${\bm{J}}({\bm{x}},y)$ w.r.t. to ${\bm{x}}$: ${\bm{x}}^{\prime}={\bm{x}}+\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime},y)),$ where the noise meets the $L^{\infty}$ norm bound $\epsilon$. Furthermore, this approach can be applied iteratively, as shown by (Kurakin et al., 2018), using a reduced step size $\alpha$: ${\bm{x}}^{\prime}_{0}={\bm{x}},\;{\bm{x}}^{\prime}_{t+1}={\bm{x}}^{\prime}_{t}+\alpha\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime}_{t},y))$, where in each step, the perturbation should be constrained within the $L^{\infty}$ ball of radius $\epsilon$. This constraint is a characteristic of the Projected Gradient Descent (PGD) attack proposed by (Madry et al., 2017), which is commonly considered a standard attack for the evaluation of model robustness (Liu et al., 2023a). Masked PGD (mPGD) (Xu et al., 2023b) is a variant of the PGD attack that restricts perturbations to a specific region within an image. The PGD attack is extended towards a patch: ${\bm{x}}_{t+1}^{\prime}=\text{Clip}_{{\bm{x}}^{\prime}}\left({\bm{x}}_{t}^{\prime}+\alpha\cdot\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}_{t}^{\prime},y,\bm{\theta})[patch]),\epsilon\right).$ In this context, the term “patch” refers to the region specified as $[x:x+h,y:y+w]$, where $[x,y,h,w]$ are the provided patch’s coordinates and dimensions.
Lastly, DeepFool (DF) (Moosavi-Dezfooli et al., 2016) assumes that the network’s decision boundary is linear, even though in reality, it may not be. It aims to find the minimal perturbation, corresponding to the orthogonal projection onto the hyperplane.
2.2 Advantages of Post-Hoc OOD Detectors
Post-hoc OOD detection methods, which make use of specific layers of the pre-trained classifier, have been demonstrated to outperform retraining-based approaches, thereby underscoring their empirical efficacy (Zhang et al., 2023). Their plug-and-play nature allows seamless integration with pre-trained models, without necessitating alterations to the training procedure or access to original training data (Zhang et al., 2023). In Table 1 are 16 post-hoc detectors listed, which also shows if a method uses features, logits, or the probabilities of the pre-trained model. Post-hoc methods underline their simplicity to outperform others on natural distribution shifts datasets by being as lightweight as possible. The latest post-hoc OOD detector in Table 1 is SCALE (Xu et al., 2023a). In contrast to the previous activation shaping method, ASH (Djurisic et al., 2022), that involves pruning and scaling of activations, SCALE demonstrates state-of-the-art results are solely derived from scaling.
OOD detection has a more expansive scope when compared to anomaly detection or open-set recognition (OSR) (Scheirer et al., 2012). Anomaly detection is concerned with the identification of rare deviations within a single distribution. OSR addresses the issue of unknown classes during inference. OOD detection methods aim to identify any test sample that deviates from the training data distribution (Zhang et al., 2023).
Moreover, post-hoc OOD methods can be augmented with other techniques, such as those employed in OSR (Gillert & von Lukas, 2021) or uncertainty estimation (Schwaiger et al., 2020). Combining different techniques means also the post-hoc methods become more complex and thus the attack surface might become larger, i.e. attacks against uncertainty estimation (Ledda et al., 2023) differ from the evasion attacks. It is not necessarily the case that a post-hoc method must be combined with other techniques. This demonstrates SAFE (Wilson et al., 2023).
2.3 OOD Adversarial Detection
Ensuring the protection of deployed DL models is the aim of OOD detectors, but the task of providing comprehensive defense (Carlini et al., 2019) against unknown threats is challenging. Every defense mechanism can be circumvented at some point (Carlini & Wagner, 2017a). Many OOD detectors can be easily evaded by slightly perturbing benign OOD inputs, creating OOD adversarial examples that reveal a severe limitation of current open-world learning frameworks (Sehwag et al., 2019b; Azizmalayeri et al., 2022). Even adversarial training-based defense methods, effective against ID adversarial attacks, struggle against OOD adversarial examples (Azizmalayeri et al., 2022). In the past years, various defensive techniques (Wu et al., 2023) and combinations of them have surfaced to combat threats such as adversarial training (Madry et al., 2017; Wang et al., 2023; Bai et al., 2024), gradient masking for obfuscation (Papernot et al., 2017), input transformations such as input purification (Nie et al., 2022; Lin et al., 2024). However, attackers consistently adapt their adversarial attacks to the specific defense mechanisms (Tramer et al., 2020). According to (Croce et al., 2022), optimization-based defenses could be a promising future, because they can adapt during test-time towards the input.
OOD detectors have benefited from insights in the adversarial machine learning (AML) field, but still lack comprehensive defense against unknown threats. There are adversarial training based methods, e.g. ALOE (Chen et al., 2020), OSAD (Shao et al., 2020), or ATOM (Chen et al., 2021). A discriminator-based method, ADT (Azizmalayeri et al., 2022), significantly outperforms previous methods by addressing their vulnerabilities to strong adversarial attacks. More recently, the post-hoc method SAFE (Wilson et al., 2023) leverages the most sensitive layers in a pre-trained classifier through targeted input-level adversarial perturbations. To this end, “adversarial robust” OOD detection methods lag, being a comprehensive defense against unknown and adaptive threats remains an intricate challenge.
3 ROBUSTNESS DEFINITION
The robustness definition in the field of OOD detectors has been ambiguous when it comes to attack methods. There are two categories of adversarial examples. The first one merely attacks the underlying pre-trained classifier and the second one aims to fool the OOD detector itself. Adversarial robustness can be considered in a classifier (Unified Robustness) or the OOD detector (Robust OOD Detection) according to (Karunanayake et al., 2024). In this work, we focus on unified robustness, which belong to the covariate shift. For image classification, a dataset $\mathcal{D}=\{({\bm{x}}_{i},y_{i});{\bm{x}}_{i}\in{\mathcal{X}},y_{i}\in{\mathcal{Y}}\}$ sample from a training distribution $\hat{P}_{\rm{data}}({\bm{x}},y)$ is used to train some classifier $C:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. In real-world deployments, distribution shift occurs when classifier $C$ receives data from test distribution $\hat{P}_{\rm{test}}({\bm{x}},y)$ where $\hat{P}_{\rm{data}}({\bm{x}},y)\neq\hat{P}_{\rm{test}}({\bm{x}},y)$ (Moreno-Torres et al., 2012). An OOD detector is a scoring function $s$ that maps an image ${\bm{x}}$ to a real number $\mathbb{R}$ such that some threshold $\tau$ arrives at the detection rule $f({\bm{x}}):\text{ID if }s({\bm{x}})\geq\tau\text{, OOD otherwise}$. Table 4 in Appendix A gives an overview of several detectors as well as considered ID and OOD datasets and model architectures. The ImageNet-1K dataset together with ResNet-50 architecture has become standard for ID. Popular OD datasets are iNaturalist, SUN, Places, and Textures. Some OOD detectors, such as ALOE, OSAD, ADT, ATOM (see Section 2.3), aim to be adversarial robust. They usually take evading attacks (see Section 2.1) such as FGSM or PGD as OOD. These computationally expensive methods only show empirical results on the small-scale CIFAR-10. A robust OOD detector is built to distinguish whether a perturbed input is OOD. Standardized OOD benchmark frameworks, i.e. OpenOOD (Zhang et al., 2023) or RoboDepth (Kong et al., 2024) do not include unified robustness in their benchmarks at the moment. Consequently, both frameworks, give a false sense of encompassing open-world capabilities. They focus on natural distribution (Hendrycks et al., 2021), where OOD detection in large-scale semantic space has attracted increasing attention (Hendrycks et al., 2019), see Appendix A. Some OOD datasets have issues, where ID classes are part of the OD dataset (Bitterwolf et al., 2023). Recently, (Yang et al., 2023) found a clean semantic shift dataset that minimizes the interference of covariate shift. The experiments show that state-of-the-art OOD detectors are more sensitive to covariate shift and the advances in semantic shift detection are minimal. The investigation of adversarial examples could gain insights into understanding the covariate shift and towards a generalized OOD detection (Yang et al., 2021). The difference between benign ${\bm{x}}$ and the attacked counterpart ${\bm{x}}^{\prime}$ is the different attention of the pre-trained classifier $C$ per sample. Let’s define an attention map ${\bm{A}}$ (Guo et al., 2022): The input image is passed through a classifier $C$ to obtain a feature map ${\bm{F}}$. A possible tool could be Grad-CAM (Selvaraju et al., 2017) to investigate the attention change between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ on ${\bm{F}}$ (Rieger & Hansen, 2020).
4 EXPERIMENTS
Experiment Setup
We extend the OpenOOD framework (Zhang et al., 2023) to consider adversarial attacks. We attack the pre-trained classifiers on the corresponding test sets and evaluate 16 post-hoc OOD detectors. As attack methods, we choose FGSM(-$L^{\infty}$), PGD(-$L^{\infty}$), DF(-$L^{2}$) from FoolBox (Rauber et al., 2017); and the mPGD-($L^{\infty}$). The attacked models are ResNet-18 (He et al., 2016a), ResNet-50 (He et al., 2016a), and Swin-T (Liu et al., 2021). The datasets are CIFAR-10 & CIFAR-100 (Krizhevsky et al., 2009), ImageNet-1K (Deng et al., 2009) and its variant ImageNet-200 with just 200 classes.
Table 2: Setup. The attack success rate (ASR) from various attacks on different models and datasets. $\mathcal{A}_{\text{std}}$ refers to the standard accuracy of the pre-trained classifier.
Dataset Arch $\mathcal{A}_{\text{std}}$ (%) Attack ASR (%)
CIFAR-10 ResNet-18 95.32 PGD 99.88
FGSM 59.21
DF 100
mPGD 68.06
CIFAR-100 ResNet-18 77.19 PGD 100
FGSM 91.99
DF 100
mPGD 88.14
ImageNet-200 ResNet-18 86.27 PGD 99.9
FGSM 95.46
DF 100
mPGD 96.53
ImageNet-1K ResNet-50 76.19 PGD 99.97
FGSM 93.33
DF 100
mPGD 98.48
Swin-T 95.99 PGD 99.99
FGSM 75.09
DF 100
mPGD 98.84
The efficacy of the attacks is not absolute and depends on a multitude of factors, including the hyperparameters, model architecture, and the dataset. The attack success rate (ASR) is presented in Table 2. The attacks PGD and FGSM do have an epsilon size, we use an epsilon size of $8/255$ for CIFAR-10/100 and $4/255$ for the ImageNet. The mPGD randomly attacks an area of the image ($8\times 8$ px for CIFAR-10/100 and $60\times 60$ px for the ImageNet) without an epsilon constraint, leading to perceptible perturbations.
[…]
Detector Attacks CIFAR-10 CIFAR-100 ImageNet-200 ImageNet-1K
ResNet-18 ResNet-50 Swin-T
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
SCALE PGD 99.67 34.53 99.97 16.18 95.49 35.14 100.00 0.20 95.49 35.14
FGSM 85.74 77.50 49.69 85.64 79.75 76.88 89.75 66.28 79.75 76.88
DF 67.07 81.73 69.22 68.69 79.75 76.88 87.82 57.77 79.75 76.88
mPGD 88.50 70.69 85.58 59.67 93.24 42.17 100.00 6.90 93.24 42.17
NNGUIDE PGD 99.39 30.29 98.85 17.07 96.44 33.60 100.00 0.12 96.44 33.60
FGSM 93.10 53.01 68.14 77.62 83.21 75.13 85.27 73.53 83.21 75.13
DF 92.08 63.25 85.36 64.14 83.21 75.13 82.19 62.81 83.21 75.13
mPGD 92.94 58.90 90.98 57.26 94.30 42.87 99.99 9.52 94.30 42.87
GEN PGD 99.51 41.75 99.90 26.03 89.17 40.03 100.00 0.21 89.17 40.03
FGSM 70.14 81.29 45.66 87.10 72.06 79.00 83.63 73.28 72.06 79.00
DF 44.32 85.98 71.38 65.82 72.06 79.00 80.01 62.96 72.06 79.00
mPGD 83.65 74.35 75.59 66.56 88.10 47.17 99.96 12.40 88.10 47.17
ASH PGD 99.67 31.23 99.96 24.49 97.06 32.97 100.00 0.19 97.06 32.97
FGSM 86.94 70.60 42.86 88.14 83.81 74.57 85.61 69.99 83.81 74.57
DF 77.10 74.62 74.98 65.44 83.81 74.57 83.29 60.88 83.81 74.57
mPGD 90.75 64.35 78.87 66.61 95.20 40.17 100.00 8.49 95.20 40.17
DICE PGD 96.65 36.09 99.93 23.45 95.03 34.24 100.00 0.11 95.03 34.24
FGSM 75.86 72.65 46.27 86.51 75.62 78.73 86.99 71.20 75.62 78.73
DF 68.84 73.44 76.73 65.46 75.62 78.73 84.72 62.92 75.62 78.73
mPGD 89.46 66.25 80.48 66.68 91.60 42.99 99.99 8.63 91.60 42.99
KNN PGD 64.91 69.18 90.06 43.07 85.23 55.53 78.63 55.74 85.23 55.53
FGSM 61.23 82.08 47.81 84.69 76.88 73.23 75.89 68.43 76.88 73.23
DF 38.78 85.84 78.54 63.06 76.88 73.23 86.65 58.09 76.88 73.23
mPGD 76.02 75.02 78.93 64.03 88.63 54.44 87.09 48.40 88.63 54.44
VIM PGD 92.45 56.83 98.16 42.79 89.17 46.44 100.00 5.16 89.17 46.44
FGSM 54.89 84.43 54.72 74.75 71.60 69.44 75.28 71.31 71.60 69.44
DF 43.92 84.92 78.57 61.75 71.60 69.44 82.95 59.88 71.60 69.44
mPGD 80.70 74.35 80.56 60.01 90.16 50.71 99.86 24.51 90.16 50.71
KLM PGD 91.43 60.91 91.94 45.75 90.87 54.58 95.49 40.77 90.87 54.58
FGSM 96.90 66.02 72.31 80.83 80.71 74.44 80.54 71.56 80.71 74.44
DF 80.84 71.64 91.38 59.58 80.71 74.44 85.48 59.30 80.71 74.44
mPGD 97.52 62.26 89.69 61.30 91.96 55.14 94.81 41.19 91.96 55.14
MLS PGD 99.58 39.65 99.96 24.43 94.43 35.64 100.00 0.12 94.43 35.64
FGSM 75.61 80.97 43.04 87.63 74.47 79.06 85.30 74.16 74.47 79.06
DF 51.12 84.89 74.91 65.41 74.47 79.06 81.44 63.37 74.47 79.06
mPGD 85.11 73.78 78.81 66.41 90.90 44.19 99.97 10.65 90.90 44.19
REACT PGD 98.84 45.19 99.89 25.13 94.49 35.90 100.00 4.15 94.49 35.90
FGSM 79.55 79.84 42.83 88.27 76.12 78.15 80.31 74.14 76.12 78.15
DF 54.80 84.16 74.99 65.49 76.12 78.15 79.97 62.88 76.12 78.15
mPGD 85.32 73.19 78.70 66.52 91.26 44.59 99.71 20.13 91.26 44.59
GRAM PGD 99.82 22.50 99.94 17.12 98.75 25.92 100.00 0.07 98.75 25.92
FGSM 94.77 56.54 79.42 79.23 91.60 69.07 96.06 59.67 91.60 69.07
DF 88.14 60.87 90.78 58.04 91.60 69.07 93.66 55.04 91.60 69.07
mPGD 93.21 55.56 91.49 58.72 97.83 34.92 100.00 4.36 97.83 34.92
RMDS PGD 49.03 82.70 66.28 77.08 53.69 76.40 37.47 95.23 53.69 76.40
FGSM 68.03 80.66 76.00 79.46 67.10 76.26 71.65 73.33 67.10 76.26
DF 43.26 85.75 64.52 80.96 67.10 76.26 73.74 65.11 67.10 76.26
mPGD 77.05 75.84 84.46 74.74 79.37 63.99 90.78 50.99 79.37 63.99
EBO PGD 99.58 39.61 99.96 24.49 94.47 35.41 100.00 0.12 94.47 35.41
FGSM 75.62 81.04 42.86 88.14 74.58 79.17 85.35 74.42 74.58 79.17
DF 51.23 84.71 74.99 65.44 74.58 79.17 81.63 63.79 74.58 79.17
mPGD 85.11 73.80 78.87 66.61 91.00 44.10 99.97 10.64 91.00 44.10
MDS PGD 47.81 84.48 50.82 84.18 57.34 76.11 0.05 99.95 57.34 76.11
FGSM 64.24 79.22 86.06 51.31 91.41 49.61 90.61 52.93 91.41 49.61
DF 54.04 79.60 89.76 52.53 91.41 49.61 93.19 49.18 91.41 49.61
mPGD 78.25 69.86 91.87 50.63 83.81 65.37 51.74 89.98 83.81 65.37
ODIN PGD 99.73 33.25 99.97 17.58 97.20 31.19 100.00 0.91 97.20 31.19
FGSM 73.56 83.16 39.74 90.19 77.39 76.38 85.44 72.56 77.39 76.38
DF 60.16 83.99 72.94 68.00 77.39 76.38 82.00 65.46 77.39 76.38
mPGD 86.49 74.11 85.66 62.09 91.72 45.12 99.90 15.89 91.72 45.12
MSP PGD 99.34 43.62 100.00 27.04 87.66 42.60 100.00 4.08 87.66 42.60
FGSM 67.16 81.14 47.87 85.08 72.12 77.77 80.39 70.45 72.12 77.77
DF 39.78 86.85 70.63 65.92 72.12 77.77 72.65 60.76 72.12 77.77
mPGD 82.76 74.68 75.03 66.23 87.48 49.07 100.00 22.21 87.48 49.07
4.1 DISCUSSION
All methods do not show sufficient results, when evaluating 16 post-hoc methods as unveiled in the Table 3. Only two of them - Mahalanobis distance-based methods - (Lee et al., 2018; Ren et al., 2021) show partly detection capabilities against the FGSM, PGD, and mPGD attacks on ResNet-50 with ImageNet-1K. The Mahalanobis distance’s robustness has been thoroughly studied (Kamoi & Kobayashi, 2020; Eustratiadis et al., 2021; Yang et al., 2022b; Anthony & Kamnitsas, 2023), and its adversarial robustness is attributed to its covariance structure consideration (Eustratiadis et al., 2021). This highlights the conflict between AML and OOD detection, where detectors often excel in either adversarial or natural distributions, but not both. High detection rates on OOD samples are the fundament for further defense mechanisms.
Level of Adversarial Robustness - From Detectors towards Defenses A defense is more sophisticated than detectors to mitigate the attacker’s efforts to fool a classifier. A step towards adversarial defense could be to improve adversarial robustness in OOD detectors. We suggest a possible roadmap to evaluate detectors and lift them toward an adversarial defense:
  1. 1.
    Evaluate on strong attacks (Carlini & Wagner, 2017a) and avoid hyperparameters that weaken the strength of the attack’s effect. The FGSM is not recommended because it performs a single step to find the adversarial perturbation, making it less effective than PGD (Li et al., 2020). Furthermore, the attack hyperparameter space is huge (Cinà et al., 2024) and could mitigate the attacks’ strength.
  2. 2.
    Use different models and other datasets than the simple ResNet-18 trained on CIFAR-10. We suggest using ImageNet-1K because its complexity in terms of resolution and objects is more closely to real-world scenarios.
  3. 3.
    Elaborate your strategy to countermeasure the attack as demonstrated in (Sehwag et al., 2019b). New defense mechanisms have often been broken quickly again (Carlini & Wagner, 2017a). For example, a differentiable OOD detector can be easily fooled if the attacker approximates the gradients of the network during the backward pass in a differentiable manner, known as BPDA (Athalye et al., 2018).
  4. 4.
    Let your method fail against sophisticated attacks, such as adaptive attacks (Athalye et al., 2018; Croce et al., 2022) or design OOD adversarial examples that convert OD to ID samples (Sehwag et al., 2019a). Adversarial robustness is an iterative process, where defenses are proposed, evaluated, and then improved upon in response to new attacks or discovered vulnerabilities.
5 CONCLUSION
In this study, we assess the performance of the 16 post-hoc OOD detectors in their ability to detect various evasive attacks. We conducted prominent white-box adversarial attacks, such as PGD and DeepFool, on the CIFAR-10 and ImageNet-1K datasets. Our discovery indicates that current post-hoc methods are not ready for real-world applications as long as they are vulnerable against well-known threat — adversarial examples. We hope that our experiments give a baseline for further research by improving post-hoc methods towards robustness and will find a place as a standardized benchmark, such as OpenOOD.
Future Work We propose to extend the experiments towards transferability because adversarial examples transfer effectively across different datasets (Alhamoud et al., 2022) and models (Gu et al., 2023). Finally, we would suggest using black-box attacks for a realistic open-world scenario.
Appendix A OOD Definition per Method
This section extends the Section 3. In Table 4, we compare the chosen ID and OOD datasets of each OOD detector. In this comparison, we pick those experiments with largest dimensioned available datasets. Furthermore, we extend this comparison by appending the adversarial robust OOD detectors from the related work in Section 2.3. It can be observed that the more recent post-hoc detectors tend to have ImageNet-1K as standard ID datasets. In contrast, the adversarial robust OOD detectors are benchmarked more likely on the smaller and less complex CIFAR-10 dataset.
Following attacks, models, and datasets are used to heuristically evaluate OOD samples (more details in Table 4):
Attacks: FGSM (Goodfellow et al., 2014), PGD (Madry et al., 2017)
Model Architectures:
  • ResNet-18/ResNet-50 (He et al., 2016a), ResNetv2-101 (He et al., 2016b), ResNet-50-D (Touvron et al., 2019), TResNet-M (Ridnik et al., 2021b), WideResNet (Zagoruyko & Komodakis, 2016)
  • BiT (Kolesnikov et al., 2020), VIT-B-16 (Dosovitskiy et al., 2020), ViT (Touvron et al., 2021), DeiT (Touvron et al., 2021), Swin-T (Liu et al., 2021)
  • DenseNet-121 & DenseNet-101 (Huang et al., 2017)
  • MobileNet (Howard et al., 2017), MobileNetV2 (Sandler et al., 2018)
  • RegNet & RegNetX4.0 (Radosavovic et al., 2020)
  • RepVGG (Ding et al., 2021)
  • Mixer-B-16 (Tolstikhin et al., 2021)
  • CLIP (Radford et al., 2021)
Datasets: CIFAR-10/CIFAR-100 (Krizhevsky et al., 2009), BDD-Anomaly (Hendrycks et al., 2019), DeepDrive (Wu et al., 2022) ImageNet-1K (Deng et al., 2009), ImageNet-21K (Ridnik et al., 2021a), ImageNet-O (Hendrycks & Dietterich, 2019), iNaturalist (Van Horn et al., 2018), ISUN (Quattoni & Torralba, 2009), GSTRB (Stallkamp et al., 2012), LSUN (Yu et al., 2015), NINCO (Bitterwolf et al., 2023), OpenImage-O (Wang et al., 2022), Pascal-VOC (Hoiem et al., 2009), Places (Zhou et al., 2017), Textures (Cimpoi et al., 2014), TinyImageNet (Le & Yang, 2015), Species (Van Horn et al., 2015), StreetHazards (Hendrycks et al., 2019), SSB-hard (Vaze et al., 2021), SVHN (Netzer et al., 2011), SUN (Xiao et al., 2010).
Table 4: Overview of ID and OOD definition from several OOD detectors. The detectors are divided into the following categories (Zhang et al., 2023): Classifiction-based, Density-based, Distance-based. We also mark: Supervised and Adversarial Robust.
Methods ID OOD Model Architectures
PostHoc Methods
SCALE (Xu et al., 2023a) ImageNet-1K Near-OOD: NINCO, SSB-hard; Far-OOD: iNaturalist, OpenImage-O, Textures ResNet-50
NNGuide (Park et al., 2023) ImageNet-1K Near-OOD: iNaturalist, OpenImage-O; Far-OOD: Textures; Overlapping: SUN and Places MobileNet, RegNet, ResNet-50, ViT
GEN (Liu et al., 2023b) ImageNet-1K ImageNet-O, iNaturalist, OpenImage-O, Texture, BiT, DeiT, RepVGG, ResNet-50, ResNet-50-D, Swin-T, ViT
ASH (Djurisic et al., 2022) ImageNet-1K iNaturalist, Places, SUN, Textures MobileNetV2, ResNet-50
DICE (Sun & Li, 2022) ImageNet-1K iNaturalist, Places, SUN, Textures DenseNet-101
KNN (Sun et al., 2022) ImageNet-1K iNaturalist, Places, SUN, Textures ResNet-50
VIM (Wang et al., 2022) ImageNet-1K ImageNet-O, iNaturalist, OpenImage-O, Texture BiT-S, DeiT, RepVGG, ResNet-50, ResNet-50-D, Swin-T, VIT-B-16
KLM; MLS (Hendrycks et al., 2019) ImageNet-21K; ImageNet-1K, Places Species (categories); BDD-Anomaly, StreetHazards (segmentation) Mixer-B-16; ResNet-50, TResNet-M, ViTB-16
REACT (Sun et al., 2021) ImageNet-1K iNaturalist, Places, SUN, Textures MobileNet, ResNet
GRAM (Huang et al., 2021) ImageNet-1K iNaturalist, SUN, Places, Textures DenseNet-121, ResNetv2-101
RMDS (Ren et al., 2021) CIFAR-10, CIFAR-100 CIFAR-10, CIFAR-100 BiT, CLIP, VIT-B-16
EBO (Liu et al., 2020) CIFAR-10 ISUN, Places, Texture, SVHN, LSUN WideResNet
MDS (Lee et al., 2018) CIFAR-10 SVHN, TinyImageNet, LSUN, Adversarial Examples DenseNet, ResNet
ODIN (Liang et al., 2017) CIFAR-10 LSUN, SVHN, TinyImageNet DenseNet, ResNet
MSP (Hendrycks & Gimpel, 2016) CIFAR-10 SUN (Gaussian) WideResNet 40-4
OOD Detectors for Adversarial Robustness
ALOE (Chen et al., 2020) CIFAR-10, CIFAR-100, GSTRB PGD attack DenseNet
OSAD (Shao et al., 2020) CIFAR-10, SVHN, TinyImageNet FGSM, PGD attack ResNet-18
ADT (Azizmalayeri et al., 2022) CIFAR-10, CIFAR-100 FGSM, PGD attack ViT
ATOM (Chen et al., 2021) CIFAR-10, CIFAR-100, SVHN PGD attack WideResNet
SAFE (Wilson et al., 2023) PASCAL-VOC, DeepDrive FGSM attack RegNetX4.0, ResNet-50