### v3

# DECIPHERING THE DEFINITION OF ADVERSARIAL ROBUSTNESS FOR POST-HOC OOD DETECTORS

Abstract
Detecting out-of-distribution (OOD) inputs is critical for safely deploying deep learning models in real-world scenarios. In recent years, many OOD detectors have been developed, and even the benchmarking has been standardized, i.e. OpenOOD. The number of post-hoc detectors is growing fast and showing an option to protect a pre-trained classifier against natural distribution shifts, claiming to be ready for real-world scenarios. However, its efficacy in handling adversarial examples has been neglected in the majority of studies. This paper investigates the adversarial robustness of 16 post-hoc detectors on several evasion attacks and discusses a roadmap toward adversarial defense in OOD detectors.
Keywords adversarial examples  $\cdot$ OOD  $\cdot$ post-hoc detectors
1 INTRODUCTION
Adversarial robustness in the context of out-of-distribution (OOD) detection refers to the ability of a detector to correctly identify OOD samples even when they have been adversarial perturbed to evade a deep neural network (DNN). Evasion attacks, e.g. PGD [1], which are designed to fool deep learning classifiers, are difficult to spot as an outlier for OOD detectors. To prevent errors in real-world applications, it is crucial to detect OOD cases, not only for natural distribution shifts [2] but also for adversarial examples [3] without degrading the generalizability of the underlying pre-trained classifier [4].
Current standardized benchmarks such as OpenOOD [5] and RoboDepth [6] merely focus on natural distribution shifts and corruptions [7]. Especially OpenOOD aims to make a fair comparison across methods initially developed for anomaly detection [8], model uncertainty [9], open-set recognition (OSR) [10], and OOD detection [11] but neglecting adversarial examples.
OpenOOD benchmark suite [4] evaluates methods on semantic shift (e.g., samples that are semantically different from the training data, representing truly novel or unseen concepts.) [12] and covariate shift (e.g. samples that come from a different distribution than the training data, but still belong to the same semantic categories).
[…]
Detector Venue Detector Architecture Adversarial
Features Logits Probs Robust
SCALE ICLR’24 $\checkmark$ $\checkmark$
NNGUIDE NeurIPS’23 $\checkmark$ $\checkmark$
GEN CVPR’23 $\checkmark$
ASH ICLR’23 $\checkmark$
DICE ECCV’22 $\checkmark$
KNN ICML’22 $\checkmark$
VIM CVPR’22 $\checkmark$ $\checkmark$
KLM ICML’22 $\checkmark$
MLS ICML’22 $\checkmark$
REACT NeurIPS’21 $\checkmark$
[…]
GRAM ICML’20 $\checkmark$
EBO NeurIPS’20 $\checkmark$
ODIN ICLR’18 $\checkmark$
[…]
MSP ICLR’17 $\checkmark$
Current OOD detection methods, as some listed in Table 1, achieve outstanding results on prominent OOD benchmarks, such as the OpenImage-O [13], ImageNet-O [14], Texture [15], and iNaturalist [16, 17]. OOD detection is a very quickly growing field by the number of methods added to OpenOOD. More specific post-hoc methods with their plug-and-play capabilities on pre-trained classifiers are more flexible and scalable compared to methods that require full retraining on new OOD data [18, 19]. Simple post-hoc methods like KNN [20] are highlighted maintaining good performance on toy-datasets (e.g. MNIST [21], CIFAR-10 or CIAR-100 [22]), and also show outstanding performance on the more realistic dataset like ImageNet [23] according to [18]. These experiments do neglect the adversarial robustness of “state-of-the-art” (sota) detectors and the real-world capabilities [24] are questionable as past studies had shown [3, 25, 26, 27]. OOD detectors should be able to reliably detect inputs that come from a different distribution than the training data, even if the shift is subtle or complex. Adversarial examples remain challenging because they share the same semantics as the training data, but aim to modify the classifier’s output.
In this study, we investigate the adversarial robustness of post-hoc ODD detectors. Our contributions can be summarized as follows:
  • We revise the definition of adversarial OOD regarding post-hoc OOD methods to finally have a common understanding of robust adversarial OOD detection.
  • We examine 16 post-hoc OOD detectors by delving into their current ability to detect adversarial examples — an aspect that has been disregarded.
  • We expand the OpenOOD framework with evasive attacks and provide adversarial OOD datasets: github.com/adverML/AdvOpenOOD.
2 RELATED WORK
This section discusses existing white-box evasion attacks, the benefits of post-hoc detectors, existing methods for detecting adversarial samples, and finally Grad-CAM, a tool for understanding the neural networks’ attention shift initiated by adversarial examples.
2.1 Evasion Attacks Crafting Inliers
The objective of evasion attacks is to generate adversarial examples that will result in the misclassification of inputs at deep learning models [28]. It is possible to distinguish between two types of attack: black-box attacks [29], where the classifier is queried, and white-box attacks [30], where the network is under the attacker’s complete control. A white-box threat model is strictly stronger. They try to find the smallest possible perturbation, often imperceptible to humans, to manipulate the model’s decision boundaries.
More formally, for an input ${\bm{x}}$ with the ground-truth label $y$, an adversarial example ${\bm{x}}^{\prime}$ is crafted by adding small noise $\bm{\delta}$ to ${\bm{x}}$ such that the predictor model loss ${\bm{J}}({\bm{x}}^{\prime},y)$ is maximized. The $L^{p}$ norm of the adversarial noise should be less than a specified value $\epsilon$, i.e., $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, e.g., $\epsilon=8/255$ [31], to ensure that the image does not change semantically. The attack method Fast Gradient Sign Method (FGSM) by [32] maximizes the loss function in a single step by taking a step towards the sign of the gradient of ${\bm{J}}({\bm{x}},y)$ w.r.t. to ${\bm{x}}$: ${\bm{x}}^{\prime}={\bm{x}}+\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime},y))$, where the noise meets the $L^{\infty}$ norm bound $\epsilon$. Furthermore, this approach can be applied iteratively, as shown by [33], using a reduced step size $\alpha$: $vx^{\prime}_{0}={\bm{x}},\;{\bm{x}}^{\prime}_{t+1}={\bm{x}}^{\prime}_{t}+\alpha\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime}_{t},y))$, where in each step, the perturbation should be constrained within the $L^{\infty}$ ball of radius $\epsilon$. This constraint is a characteristic of the Projected Gradient Descent (PGD) attack proposed by [1], which is commonly considered a standard attack for the evaluation of model robustness [34]. Masked PGD (mPGD) [35] is a variant of the PGD attack that restricts perturbations to a specific region within an image. The PGD attack is extended towards a patch: ${\bm{x}}_{t+1}^{\prime}=\text{Clip}_{{\bm{x}}^{\prime}}\left({\bm{x}}_{t}^{\prime}+\alpha\cdot\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}_{t}^{\prime},y,\bm{\theta})[patch]),\epsilon\right).$ In this context, the term “patch” refers to the region specified as $[x:x+h,y:y+w]$, where $[x,y,h,w]$ are the provided patch’s coordinates and dimensions.
Lastly, DeepFool (DF) [36] assumes that the network’s decision boundary is linear, even though in reality, it may not be. It aims to find the minimal perturbation, corresponding to the orthogonal projection onto the hyperplane.
2.2 Advantages of Post-Hoc OOD Detectors
Post-hoc OOD detection methods, which make use of specific layers of the pre-trained classifier, have been demonstrated to outperform retraining-based approaches, thereby underscoring their empirical efficacy [5]. Their plug-and-play nature allows seamless integration with pre-trained models, without necessitating alterations to the training procedure or access to original training data [5]. In Table 1 are 16 post-hoc detectors listed, which also shows if a method uses features, logits, or the probabilities of the pre-trained model. Post-hoc methods underline their simplicity to outperform others on natural distribution shifts datasets by being as lightweight as possible. The latest post-hoc OOD detector in Table 1 is SCALE [37]. In contrast to the previous activation shaping method, ASH [38], that involves pruning and scaling of activations, SCALE demonstrates sota results are solely derived from scaling.
OOD detection has a more expansive scope when compared to anomaly detection or open-set recognition (OSR) [39]. Anomaly detection is concerned with the identification of rare deviations within a single distribution. OSR addresses the issue of unknown classes during inference. OOD detection methods aim to identify any test sample that deviates from the training data distribution [5].
Moreover, post-hoc OOD methods can be augmented with other techniques, such as those employed in OSR [40] or uncertainty estimation [41]. Combining different techniques means also the post-hoc methods become more complex and thus the attack surface might become larger, i.e. attacks against uncertainty estimation [9] differ to the evasion attacks. It is not necessarily the case that a post-hoc method must be combined with other techniques. This demonstrates SAFE [42].
2.3 OOD Adversarial Detection
Ensuring the protection of deployed DL models is the aim of OOD detectors, but the task of providing comprehensive defense [43] against unknown threats is challenging. Every defense mechanism can be circumvented at some point [44]. Many OOD detectors can be easily evaded by slightly perturbing benign OOD inputs, creating OOD adversarial examples that reveal a severe limitation of current open-world learning frameworks [45, 46]. Even adversarial training-based defense methods, effective against ID adversarial attacks, struggle against OOD adversarial examples [46]. In the past years, various defensive techniques [47] and combinations of them have surfaced to combat threats such as adversarial training [1, 48, 49], gradient masking for obfuscation [50], input transformations such as input purification [51, 52]. However, attackers consistently adapt their adversarial attacks to the specific defense mechanisms [53]. According to [54], optimization-based defenses could be a promising future, because they can adapt during test-time towards the input.
OOD detectors have benefited from insights in the adversarial machine learning (AML) field, but still lack comprehensive defense against unknown threats. There are adversarial training based methods, e.g. ALOE [26], OSAD [55], or ATOM [56]. A discriminator-based method, ADT [46], significantly outperforms previous methods by addressing their vulnerabilities to strong adversarial attacks. More recently, the post-hoc method SAFE [42] leverages the most sensitive layers in a pre-trained classifier through targeted input-level adversarial perturbations. To this end, “adversarial robust” OOD detection methods lag, being a comprehensive defense against unknown and adaptive threats remains an intricate challenge.
2.4 CAM-based Explanations and Consistent Shift by Adversarial Examples
Explainable AI (XAI) [57] can be used as a defense mechanism [58, 59, 60] to spot adversarial examples [61]. The Grad-CAM heatmap provides a clear indication, for a given image, of the regions that the Convolutional Neural Network (CNN) focuses on to make a decision. Class Activation Maps (CAMs) can be considered saliency maps specific to the input [62]. These maps emerge from the accumulated and up-scaled activations at a particular convolutional layer, typically the penultimate layer. The classification is estimated through a linear combination of the activation of units in the final layer(s) of the feature selection network:
$f_{\theta}(.)\approx\sum_{i}\sum_{k}w_{k}a_{ki},$ (1)
where $a_{ki}$ represents the activation of the k-th channel of unit $i$, and $w_{k}$ denotes the learned weights. The relevance values are subsequently defined as $r_{i}=\sum_{k}w_{k}a_{ki}$. The determination of these weights depends on the specific CAM variant employed [63, 64, 65]. Grad-CAM weights the activations using gradients:
$w_{K}:=\frac{\partial f_{\theta}(.)}{\partial a_{ki}}.$ (2)
This weighting directly links to more fundamental explanations that merely estimate the influence of the input on the final output as described before $r_{i}=\partial f_{\theta}(.)/\partial{\bm{x}}_{i}$ [66].
In context to adversarial examples, Chakraborty et al. [67] observes a consistent shift in the region highlighted in the gradient activation heatmap, reflecting its contribution to decision-making across all models under adversarial attacks.
3 DEFINITIONS IN OOD DETECTION
Table 2: Overview of ID and OOD definition from several OOD detectors. The post-hoc detectors are divided into the following categories [5]: Classifiction-based, Density-based, Distance-based. We also mark: Supervised and Adversarial Robust, which are not post-hoc detectors.
Methods ID OOD Model Architectures
Post-Hoc OOD Detection Methods
SCALE [37] ImageNet-1K Near-OOD: NINCO, SSB-hard; Far-OOD: iNaturalist, OpenImage-O, Textures ResNet-50
NNGuide [68] ImageNet-1K Near-OOD: iNaturalist, OpenImage-O; Far-OOD: Textures; Overlapping: SUN and Places MobileNet, RegNet, ResNet-50, ViT
GEN [69] ImageNet-1K ImageNet-O, iNaturalist, OpenImage-O, Texture, BiT, DeiT, RepVGG, ResNet-50, ResNet-50-D, Swin-T, ViT
ASH [38] ImageNet-1K iNaturalist, Places, SUN, Textures MobileNetV2, ResNet-50
DICE [70] ImageNet-1K iNaturalist, Places, SUN, Textures DenseNet-101
KNN [20] ImageNet-1K iNaturalist, Places, SUN, Textures ResNet-50
VIM [13] ImageNet-1K ImageNet-O, iNaturalist, OpenImage-O, Texture BiT-S, DeiT, RepVGG, ResNet-50, ResNet-50-D, Swin-T, VIT-B-16
KLM; MLS [71] ImageNet-21K; ImageNet-1K, Places Species (categories); BDD-Anomaly, StreetHazards (segmentation) Mixer-B-16; ResNet-50, TResNet-M, ViTB-16
REACT [72] ImageNet-1K iNaturalist, Places, SUN, Textures MobileNet, ResNet
GRAM [73] ImageNet-1K iNaturalist, SUN, Places, Textures DenseNet-121, ResNetv2-101
RMDS [74] CIFAR-10, CIFAR-100 CIFAR-10, CIFAR-100 BiT, CLIP, VIT-B-16
EBO [75] CIFAR-10 ISUN, Places, Texture, SVHN, LSUN WideResNet
MDS [11] CIFAR-10 SVHN, TinyImageNet, LSUN, Adversarial Examples DenseNet, ResNet
ODIN [76] CIFAR-10 LSUN, SVHN, TinyImageNet DenseNet, ResNet
MSP [12] CIFAR-10 SUN (Gaussian) WideResNet 40-4
Adversarial Robust OOD Detectors
ALOE [26] CIFAR-10, CIFAR-100, GSTRB PGD attack DenseNet
OSAD [55] CIFAR-10, SVHN, TinyImageNet FGSM, PGD attack ResNet-18
ADT [46] CIFAR-10, CIFAR-100 FGSM, PGD attack ViT
ATOM [56] CIFAR-10, CIFAR-100, SVHN PGD attack WideResNet
SAFE [42] PASCAL-VOC, DeepDrive FGSM attack RegNetX4.0, ResNet-50
The robustness definition in the field of OOD detectors has been ambiguous when it comes to attack methods. There are two categories of adversarial examples. The first one merely attacks the underlying pre-trained classifier and the second one aims to fool the OOD detector itself. Adversarial robustness can be considered in a classifier (Unified Robustness) or the OOD detector (Robust OOD Detection) according to [77]. This work focuses on unified robustness, which belongs to the covariate shift. For image classification, a dataset $\mathcal{D}=\{({\bm{x}}_{i},y_{i});{\bm{x}}_{i}\in{\mathcal{X}},y_{i}\in{\mathcal{Y}}\}$ sample from a training distribution $\hat{P}_{\rm{data}}({\bm{x}},y)$ is used to train some classifier $C:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. In real-world deployments, distribution shift occurs when classifier $C$ receives data from test distribution $\hat{P}_{\rm{test}}({\bm{x}},y)$ where $\hat{P}_{\rm{data}}({\bm{x}},y)\neq\hat{P}_{\rm{test}}({\bm{x}},y)$ [78]. An OOD detector is a scoring function $s$ that maps an image ${\bm{x}}$ to a real number $\mathbb{R}$ such that some threshold $\tau$ arrives at the detection rule $f({\bm{x}}):\text{ID if }s({\bm{x}})\geq\tau\text{, OOD otherwise}$.
Existing Definitions
In Table 2, we present a comparison of the selected ID and OOD datasets utilized by each OOD detector. Our focus lies on experiments employing the most extensive available datasets. Additionally, we enrich this analysis by including adversarial robust OOD detectors discussed in Section 2.3. Notably, recent post-hoc detectors predominantly utilize ImageNet-1K as their standard ID datasets, whereas adversarial robust OOD detectors are more commonly evaluated on the smaller CIFAR-10 dataset. Furthermore, the architecture of adversarial robust methods diverges from lightweight post-hoc designs, often relying on adversarial training and auxiliary data, as elaborated in Section 2.3. The popular adversarial ODD are the FGSM and PGD attacks. Table 2 gives an overview of several detectors as well as considered ID and OOD datasets and model architectures. The ImageNet-1K dataset together with ResNet-50 architecture has become standard for ID. Popular OD datasets are iNaturalist, SUN, Places, and Textures. Some OOD detectors, such as ALOE, OSAD, ADT, and ATOM (see Section 2.3), aim to be robust adversarial. They usually take evading attacks (see Section 2.1) such as FGSM or PGD as OOD. These computationally expensive methods only show empirical results on the small-scale CIFAR-10. A robust OOD detector is built to distinguish whether a perturbed input is OOD. Standardized OOD benchmark frameworks, i.e. OpenOOD [5] or RoboDepth [6] do not currently include unified robustness in their benchmarks. Consequently, both frameworks, give a false sense of encompassing open-world capabilities. They focus on natural distribution [14], where OOD detection in large-scale semantic space has attracted increasing attention [71]. Some OOD datasets have issues, where ID classes are part of the OD dataset [79]. Recently, [80] found a clean semantic shift dataset that minimizes the interference of covariate shift. The experiments show that sota OOD detectors are more sensitive to covariate shifts, and the advances in semantic shift detection are minimal.
Extension to Adversarial Robust Definition
Investigating adversarial examples could gain insights into understanding the covariate shift and towards a generalized OOD detection [4]. The difference between benign ${\bm{x}}$ and the attacked counterpart ${\bm{x}}^{\prime}$ is the different attention of the pre-trained classifier $C$ per sample. Let us define an attention map $A$ [81]: The input image is passed through a classifier $C$ to obtain a feature map ${\bm{F}}$. A possible tool could be Grad-CAM [64] (see also Section 2.4) to investigate the attention change between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ on ${\bm{F}}$ [60] (compare Section 4.2).
Let $\Omega({\bm{x}})$ be a set of semantic-preserving perturbations on an input ${\bm{x}}$. For $\bm{\delta}\in\Omega({\bm{x}})$, ${\bm{x}}+\bm{\delta}$ have the same distributional membership (i.e., ${\bm{x}}$ and ${\bm{x}}+\bm{\delta}$ both belong to ID or OOD) [26]. To clarify, the attention maps ${\bm{A}}$ between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ are not the same
${\bm{A}}_{C}({\bm{x}})\neq{\bm{A}}_{C}({\bm{x}}^{\prime})\text{.}$ (3)
So adversarial examples are not from a completely different distribution, but rather adversarial perturbed versions of ID data.
4 EXPERIMENTS
In this section, we explain our experimental setup and discuss the results of 16 post-hoc OOD detectors on different datasets and pre-trained models.
4.1 Experiment Setup
We extend the OpenOOD framework [5] to consider adversarial attacks. We attack the pre-trained classifiers on the corresponding test sets and evaluate 16 post-hoc OOD detectors.
As attack methods, we choose FGSM(-$L^{\infty}$), PGD(-$L^{\infty}$), DF(-$L^{2}$) from FoolBox [82]; and the mPGD-($L^{\infty}$). The attacked models are ResNet-18 [83], ResNet-50 [83], and Swin-T [84].
The datasets are CIFAR-10 & CIFAR-100 [22], ImageNet-1K [23] and its variant ImageNet-200 with just 200 classes.
Table 3: Setup. The attack success rate (ASR) from various attacks on different models and datasets. The FGSM attack has the lowest, while DF has the highest ASR. $\mathcal{A}_{\text{std}}$ refers to the standard accuracy of the pre-trained classifier.
Dataset Arch $\mathcal{A}_{\text{std}}$ (%) Attack ASR (%)
CIFAR-10 ResNet-18 95.32 PGD 99.88
FGSM 59.21
DF 100
mPGD 68.06
CIFAR-100 ResNet-18 77.19 PGD 100
FGSM 91.99
DF 100
mPGD 88.14
ImageNet-200 ResNet-18 86.27 PGD 99.9
FGSM 95.46
DF 100
mPGD 96.53
ImageNet-1K ResNet-50 76.19 PGD 99.97
FGSM 93.33
DF 100
mPGD 98.48
Swin-T 95.99 PGD 99.99
FGSM 75.09
DF 100
mPGD 98.84
The efficacy of the attacks is not absolute and depends on a multitude of factors, including the hyperparameters, model architecture, and the dataset. The attack success rate (ASR) is presented in Table 3. The attacks PGD and FGSM do have an epsilon size, we use an epsilon size of $8/255$ for CIFAR-10/100 and $4/255$ for the ImageNet. The mPGD randomly attacks an area of the image ($8\times 8$ px for CIFAR-10/100 and $60\times 60$ px for the ImageNet) without an epsilon constraint, leading to perceptible perturbations.
[…]
Figure 1: Grad-CAM comparison between the benign and its attacked counterpart. The color red indicates a high intensity of samples on similar heatmaps. The color blue indicates a low intensity. The attacks can be compared row-wise that is sorted according to datasets and attacked DNN. The DF attack yields very similar heatmaps across all datasets and models.
4.2 Grad-CAM Similarity
This section investigates the neural networks’ attention variations to classify benign and adversarial images. Our goal is to measure if adversarial attacks produce a change in the attention of the neural network that leads to misclassification as depicted in Figure 1. A previous study showed that misclassification is correlated with a shift in the center of attention for the of image classification [67]. We want to test this attention change in our pair of benign/attacked images. To quantitative measure the aforementioned attention change we use the Grad-CAM method [63]. The Grad-CAM method allows us to map the relevant part of a given image that leads to the network classification. Usually, the last block of the model is used to obtain attention [85]. In our work, we investigate the Grad-CAM output of benign and attacked images. To quantitatively analyze the Grad-CAM difference of both samples we consider two well-known metrics: the mean square error distance (L2), and the structural similar index method (SSIM) [86]. The L2 distance allows us to evaluate the pixel-wise difference of the Grad-CAM of both images. The higher the L2 value, the more distinctive the overall pixel-wise attention of the network between the benign and attacked image. The SSIM metric considers the overall structural similarity between the Grad-CAM of both images. The higher the metric the more similar the overall structure of the attention visualized through Grad-CAM maps. The combination of both metrics for our quantitative analyses is convenient because it allows us to understand better how/if an adversarial attack changes the attention of the network in an image for its classification:
  • High L2 and low SSIM report total dissimilarity of the Grad-CAM maps of the benign and attacked image. Reporting that the Grad-CAM of the attacked image is different from the benign images.
  • High L2 and high SSIM report pixel-wise differences of the Grad-CAM, but overall structural similarity. This suggests that after the attack, the attention maps shift to a different area of the image, as noted in [67].
  • Low L2 and high SSIM report total similarity, structural and pixel-wise. Reporting that the Grad-CAM is not sensitive to the changes caused by the adversarial attack.
When the L2 value approaches zero, the SSIM invariably reaches one, eliminating the possibility of simultaneously having low L2 and low SSIM. In the Figure 1, the DF attack shows a very high Grad-CAM similarity across all attack methods. Despite that, there is still a high dissimilarity by a high number of samples.
[…]
Detector Attacks CIFAR-10 CIFAR-100 ImageNet-200 ImageNet-1K ResNet-18 ResNet-50 Swin-T FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC SCALE PGD 99.67 34.53 99.97 16.18 95.49 35.14 100.00 0.20 95.49 35.14 FGSM 85.74 77.50 49.69 85.64 79.75 76.88 89.75 66.28 79.75 76.88 DF 67.07 81.73 69.22 68.69 79.75 76.88 87.82 57.77 79.75 76.88 mPGD 88.50 70.69 85.58 59.67 93.24 42.17 100.00 6.90 93.24 42.17 NNGUIDE PGD 99.39 30.29 98.85 17.07 96.44 33.60 100.00 0.12 96.44 33.60 FGSM 93.10 53.01 68.14 77.62 83.21 75.13 85.27 73.53 83.21 75.13 DF 92.08 63.25 85.36 64.14 83.21 75.13 82.19 62.81 83.21 75.13 mPGD 92.94 58.90 90.98 57.26 94.30 42.87 99.99 9.52 94.30 42.87 GEN PGD 99.51 41.75 99.90 26.03 89.17 40.03 100.00 0.21 89.17 40.03 FGSM 70.14 81.29 45.66 87.10 72.06 79.00 83.63 73.28 72.06 79.00 DF 44.32 85.98 71.38 65.82 72.06 79.00 80.01 62.96 72.06 79.00 mPGD 83.65 74.35 75.59 66.56 88.10 47.17 99.96 12.40 88.10 47.17 ASH PGD 99.67 31.23 99.96 24.49 97.06 32.97 100.00 0.19 97.06 32.97 FGSM 86.94 70.60 42.86 88.14 83.81 74.57 85.61 69.99 83.81 74.57 DF 77.10 74.62 74.98 65.44 83.81 74.57 83.29 60.88 83.81 74.57 mPGD 90.75 64.35 78.87 66.61 95.20 40.17 100.00 8.49 95.20 40.17 DICE PGD 96.65 36.09 99.93 23.45 95.03 34.24 100.00 0.11 95.03 34.24 FGSM 75.86 72.65 46.27 86.51 75.62 78.73 86.99 71.20 75.62 78.73 DF 68.84 73.44 76.73 65.46 75.62 78.73 84.72 62.92 75.62 78.73 mPGD 89.46 66.25 80.48 66.68 91.60 42.99 99.99 8.63 91.60 42.99 KNN PGD 64.91 69.18 90.06 43.07 85.23 55.53 78.63 55.74 85.23 55.53 FGSM 61.23 82.08 47.81 84.69 76.88 73.23 75.89 68.43 76.88 73.23 DF 38.78 85.84 78.54 63.06 76.88 73.23 86.65 58.09 76.88 73.23 mPGD 76.02 75.02 78.93 64.03 88.63 54.44 87.09 48.40 88.63 54.44 VIM PGD 92.45 56.83 98.16 42.79 89.17 46.44 100.00 5.16 89.17 46.44 FGSM 54.89 84.43 54.72 74.75 71.60 69.44 75.28 71.31 71.60 69.44 DF 43.92 84.92 78.57 61.75 71.60 69.44 82.95 59.88 71.60 69.44 mPGD 80.70 74.35 80.56 60.01 90.16 50.71 99.86 24.51 90.16 50.71 KLM PGD 91.43 60.91 91.94 45.75 90.87 54.58 95.49 40.77 90.87 54.58 FGSM 96.90 66.02 72.31 80.83 80.71 74.44 80.54 71.56 80.71 74.44 DF 80.84 71.64 91.38 59.58 80.71 74.44 85.48 59.30 80.71 74.44 mPGD 97.52 62.26 89.69 61.30 91.96 55.14 94.81 41.19 91.96 55.14 MLS PGD 99.58 39.65 99.96 24.43 94.43 35.64 100.00 0.12 94.43 35.64 FGSM 75.61 80.97 43.04 87.63 74.47 79.06 85.30 74.16 74.47 79.06 DF 51.12 84.89 74.91 65.41 74.47 79.06 81.44 63.37 74.47 79.06 mPGD 85.11 73.78 78.81 66.41 90.90 44.19 99.97 10.65 90.90 44.19 REACT PGD 98.84 45.19 99.89 25.13 94.49 35.90 100.00 4.15 94.49 35.90 FGSM 79.55 79.84 42.83 88.27 76.12 78.15 80.31 74.14 76.12 78.15 DF 54.80 84.16 74.99 65.49 76.12 78.15 79.97 62.88 76.12 78.15 mPGD 85.32 73.19 78.70 66.52 91.26 44.59 99.71 20.13 91.26 44.59 GRAM PGD 99.82 22.50 99.94 17.12 98.75 25.92 100.00 0.07 98.75 25.92 FGSM 94.77 56.54 79.42 79.23 91.60 69.07 96.06 59.67 91.60 69.07 DF 88.14 60.87 90.78 58.04 91.60 69.07 93.66 55.04 91.60 69.07 mPGD 93.21 55.56 91.49 58.72 97.83 34.92 100.00 4.36 97.83 34.92 RMDS PGD 49.03 82.70 66.28 77.08 53.69 76.40 37.47 95.23 53.69 76.40 FGSM 68.03 80.66 76.00 79.46 67.10 76.26 71.65 73.33 67.10 76.26 DF 43.26 85.75 64.52 80.96 67.10 76.26 73.74 65.11 67.10 76.26 mPGD 77.05 75.84 84.46 74.74 79.37 63.99 90.78 50.99 79.37 63.99 EBO PGD 99.58 39.61 99.96 24.49 94.47 35.41 100.00 0.12 94.47 35.41 FGSM 75.62 81.04 42.86 88.14 74.58 79.17 85.35 74.42 74.58 79.17 DF 51.23 84.71 74.99 65.44 74.58 79.17 81.63 63.79 74.58 79.17 mPGD 85.11 73.80 78.87 66.61 91.00 44.10 99.97 10.64 91.00 44.10 MDS PGD 47.81 84.48 50.82 84.18 57.34 76.11 0.05 99.95 57.34 76.11 FGSM 64.24 79.22 86.06 51.31 91.41 49.61 90.61 52.93 91.41 49.61 DF 54.04 79.60 89.76 52.53 91.41 49.61 93.19 49.18 91.41 49.61 mPGD 78.25 69.86 91.87 50.63 83.81 65.37 51.74 89.98 83.81 65.37 ODIN PGD 99.73 33.25 99.97 17.58 97.20 31.19 100.00 0.91 97.20 31.19 FGSM 73.56 83.16 39.74 90.19 77.39 76.38 85.44 72.56 77.39 76.38 DF 60.16 83.99 72.94 68.00 77.39 76.38 82.00 65.46 77.39 76.38 mPGD 86.49 74.11 85.66 62.09 91.72 45.12 99.90 15.89 91.72 45.12 MSP PGD 99.34 43.62 100.00 27.04 87.66 42.60 100.00 4.08 87.66 42.60 FGSM 67.16 81.14 47.87 85.08 72.12 77.77 80.39 70.45 72.12 77.77 DF 39.78 86.85 70.63 65.92 72.12 77.77 72.65 60.76 72.12 77.77 mPGD 82.76 74.68 75.03 66.23 87.48 49.07 100.00 22.21 87.48 49.07
4.3 DISCUSSION
All 16 post-hoc methods evaluated in our study as shown in Table 4 demonstrate inadequate performance. Notably, only two methods based on Mahalanobis distance [11, 74] exhibit partial detection capabilities against FGSM, PGD, and mPGD attacks on ResNet-50 with ImageNet-1K. The robustness of Mahalanobis distance has been extensively explored [87, 88, 89, 90], attributing its resilience to adversarial attacks to its consideration of covariance structures [88].
Interestingly, several OOD benchmarks [91, 18, 5, 6] and a recent challenge [92] focus solely on natural distribution shifts, neglecting adversarial examples. This underscores a fundamental conflict between Adversarial Machine Learning (AML) and Out-of-Distribution (OOD) detection, where detectors excel either in adversarial or natural distributions, but rarely both. Current OOD detectors struggle to converge on both fronts.
Adversarial examples, as counterparts of In-Distribution (ID) data, significantly alter the attention of pre-trained classifiers, as shown in Figure 1 and discussed in Equation 3. These predominantly untargeted attacks aim to minimally change class predictions, resulting in dissimilarity in attention maps (see Grad-CAM Section 4.2). The DF attack exhibits minimal attention shifts across numerous samples, yet many attacked samples still demonstrate notable deviations in attention maps. Similarly, the mPGD attack induces perceptible perturbations, with slightly improved detection results compared to PGD, but overall performance remains insufficient.
Post-hoc detectors are trained on distributions similar to the underlying pre-trained model, yet studies (e.g., [93, 94]) indicate significant discrepancies between the training distribution and adversarial distributions. Achieving high detection rates on adversarial OOD samples is crucial for developing robust defense mechanisms, a goal that remains elusive in current research.
Levels of Adversarial Robustness - From Detector towards Defense
A defense is more sophisticated than detectors to mitigate attackers’ efforts to fool a classifier. A step towards adversarial defense could be to improve adversarial robustness in OOD detectors. We suggest a possible roadmap to evaluate detectors and lift them toward an adversarial defense on different levels:
  1. Level 1:
    Evaluate on strong attacks [44] and avoid hyperparameters that weaken the strength of the attack’s effect. The FGSM is not recommended because it performs a single step to find the adversarial perturbation, making it less effective than PGD [95]. Furthermore, the attack hyperparameter space is huge [96] and could mitigate the attacks’ strength.
  2. Level 2:
    Use different models and datasets than the simple ResNet-18 trained on CIFAR-10. We suggest using ImageNet-1K because its resolution and objects are more closely related to real-world scenarios.
  3. Level 3:
    Elaborate your strategy to countermeasure the attack as demonstrated in [45]. New defense mechanisms have often been broken quickly again [44]. For example, a differentiable OOD detector can be easily fooled if the attacker approximates the gradients of the network during the backward pass in a differentiable manner, known as BPDA [97].
  4. Level 4:
    Let your method fail against sophisticated attacks, such as adaptive attacks [97, 54] or design OOD adversarial examples that convert OD to ID samples [3]. Adversarial robustness is an iterative process, where defenses are proposed, evaluated, and improved upon in response to new attacks or discovered vulnerabilities.
It must be noted that not all levels are required to be evaluated. This roadmap should indicate which level of OOD detection is robust. Further research would be encouraged to do at least an adversarial robustness benchmark (level 1), even though the OOD detector is not focused on adversarial robustness.
5 CONCLUSION
In this study, we assess the performance of the 16 post-hoc OOD detectors in their ability to detect various evasive attacks. We conducted prominent white-box adversarial attacks, such as PGD and DF, on the CIFAR-10 and ImageNet-1K datasets. Our discovery indicates that current post-hoc methods are not ready for real-world applications as long as they are vulnerable against the well-known threat — adversarial examples. We hope our experiments give a baseline for further research by improving post-hoc methods towards robustness and will find a place as a standardized benchmark, such as OpenOOD.
Future Work
We propose to extend the experiments towards transferability because adversarial examples transfer effectively across different datasets [98] and models [99]. Then, we would suggest using black-box attacks for a realistic open-world scenario. Throughout this work, we assumed a perfect pre-trained classifier on a clean curated dataset. An imperfect pre-trained classifier would be the subject of future research [100].